# Approximate predictive control barrier function for discrete-time systems

### v1

@@ Abstract
We propose integrating an explicit approximation of a predictive control barrier function (PCBF) in a safety filter framework. The approximated PCBF is implicitly defined through an optimal control problem and allows guaranteeing invariance of an implicitly defined safe set as well as stability of this safe set within a larger domain of attraction. By extending existing theoretical analysis of the PCBF, we establish inherent robustness of the original algorithm and translate the guarantees to input-to-state stability of the proposed algorithm with respect to possible approximation errors, recovering the same guarantees in the absence of approximation errors. The proposed algorithm allows certifying inputs with respect to state constraint satisfaction through a single function evaluation and filtering unsafe inputs through a control barrier function based safety filter, which is independent of the time horizon of the original predictive optimisation problem, resulting in significant online computational benefits. We demonstrate the stability properties of the proposed algorithm on a linear system example as well as its use a fast safety filter for miniature race cars in simulation.
@@ 1 INTRODUCTION
Safety filters provide a modular mechanism which enables certifying, and if necessary modifying, any control input such that safety in terms of constraint satisfaction can be guaranteed. The filters are designed to be agnostic with respect to the main control performance goal and therefore aim to minimise the distance between the proposed input and the filtered input which is applied to the system. Thereby, safety filters are suitable tools to enhance the safety properties of various systems. The filters can be used to augment human agents operating systems, such as, e.g., automotive vehicles as demonstrated on miniature race cars in [1], and have also demonstrated their effectiveness in complex robotics applications, such as, e.g., walking on stepping stones with a legged robot in [2].
@@
The basic principle of safety filters consists in obtaining a control invariant set within the state constraints, such that inputs applied to the system guarantee that the system state remains within this invariant set. In order to obtain such a control invariant set, three main branches have been pursued, as discussed extensively in [3]. Hamilton-Jacobi reachability analysis, see, e.g., [4], allows computing the maximal control invariant subset to constraints, however the method suffers from poor computational scalability with respect to increasing state dimension. Control barrier functions (CBFs), see, e.g., [5] for an overview, aim to alleviate the computational restrictions, by explicitly computing, potentially smaller, control invariant sets through either expert knowledge or more scalable computational methods compared to the Hamilton-Jacobi reachability, such as semi-definite programming or sum-of-squares optimisation, see, e.g., [6]. Finally, predictive safety filters have been proposed in [7]. They employ model predictive control (MPC) techniques, see, e.g., [8], in order to implicitly define a control invariant set through an online predictive optimisation problem. These predictive methods make use of discrete time dynamics formulations and allow to easily provide recursive feasibility guarantees as well as extensions due to the well developed MPC theory to, e.g., robust treatments of disturbances, stochastic disturbances in [9] or using online data for uncertain systems to improve safety filter performance in [10].
@@
In practical systems, considering hard constraints in predictive control methods often leads to infeasibility due to unmodeled dynamics or sensor errors. Therefore, the state constraints on the system are often softened, as proposed, e.g., in [11] and used in [1], by introducing a slack variable. However, this leads in general to a loss of theoretical properties as soon as the system leaves the constraints. In [12], a constraint softening mechanism is proposed, which is able to provide theoretical guarantees when constraints are violated. Connections of this mechanism to CBFs are investigated and it is shown that by minimising a function which depends on the slack variables, a control barrier function is implicitly defined. This result implies that the implicitly defined safe set of a predictive safety filter is stable within a larger domain of attraction, such that the system is able to recover from unexpected disturbances or model mismatch. The main drawback of the method in [12] is the need to solve two predictive optimisation problems in order to guarantee stability, which results in a computationally demanding algorithm.
1.1 Contributions
In this work, we extend initial results in [13] and aim to reduce the computational overhead of the predictive optimisation based PCBF method in [12] by obtaining an explicit approximation of the implicitly defined CBF. The explicit approximation can be used in a safety filter formulation similarly as in control barrier function based safety filters, such that only a single time step needs to be predicted using a decrease constraint on the value of the control barrier function. This decrease constraint further enables the introduction of a hyperparameter, which allows tuning the rate of stability decrease in closed-loop based on desired system behaviour. In comparison, in [12] the system is required to achieve the maximal possible decrease towards the implicitly defined safe set at every time step. We further extend the theoretical results in [12], by showing a rigorous connection between the control barrier function definition in [12] and more widely used class $\mathcal{K}$ comparator function based definitions. These relations allow to show that the method in [12] benefits from inherent robustness properties, see, e.g., [14], which are crucial for deployment of predictive control methods in real world applications. By using the insights from the inherent robustness of the algorithm proposed in [12], input-to-state stability of the proposed approximation-based algorithm is established with respect to the approximation errors. This result provides a theoretical understanding of how approximation errors in the formulation potentially impact the closed-loop performance of the system. In addition, we propose a continuous extension of the PCBF, which is negative within the implicitly defined safe set. Thereby, approximation errors lead to a significant reduction in safety filter interventions when the system is safe compared to approximating the original PCBF, which is 00 within the safe set. Finally, we show how the explicit approximation can be used in order to certify inputs using a single function evaluation, resulting in large computational speedups when the main controller proposes safe inputs.
1.2 Related work
@@
Safety filters have been extensively researched and a review of state-of-the-art methods can be found in [3]. The first main branch of research consists of control barrier function based safety filters, originally proposed in [15] for continuous time systems. The work in [16] proposed discrete-time control barrier functions. However, the proposed formulation uses a barrier function which approaches infinity as the state approaches the boundary of the safe set, which is not suitable for investigating the stability of the safe set within a larger domain. In [17], a backup CBF formulation is proposed, which allows using a backup controller to enhance an invariant safe set, by computing gradients of a CBF on the closed-loop trajectory of the backup controller. Thereby, any state which reaches the safe set under the backup controller is still considered safe, and the safe set is effectively augmented. The predictive safety filters in [18] and [12] show relations to this method, where the optimiser is free to choose all control inputs in the prediction horizon, rather than selecting them with a prior fixed policy. In [19], the predictive safety filter is augmented with a stability guarantees of the origin using a Lyapunov decrease constraint. Compared to these works, we combine the advantages of the existence of inputs which stabilise a safe set and render it invariant, while using a single CBF decrease constraint rather than a full prediction horizon, reducing the computational effort. Due to the discrete-time formulation of the predictive controllers, the resulting CBF optimisation problem is a NLP rather than a QP in [17], however the prediction horizon is not actually included in the online optimisation, effectively reducing the number of constraints. The proposed formulation in [20] uses a look-ahead to determine time instances, where given a path of the system, a CBF constraint is violated. Thereby, the control actions are modified such that safety can be guaranteed. However, this method does not consider the case when the input is constrained. Finally, [21] proposes a reachability-based safety filter formulation in continuous time, which makes use of a finite lookahead time and establishes connections to model predictive safety filter methods. However, the method uses numerical approximations of the gradients of the CBF through sampling to impose the CBF decrease and it is unclear how approximation errors can impact the closed-loop performance. In comparison, we provide a rigorous analysis of how approximation errors of the learned CBF impact the closed-loop behaviour of the system.
@@
Another main branch of research for safety filters is Hamilton-Jacobi reachability analysis, see, e.g., [4]. In [22], connections between viscosity solution of the Hamilton-Jacobi-Isaacs variational inequality, retrieving the maximum robust control invariant subset to given constraints and CBFs are established and it is shown how this solution can be used in a quadratic program. This result is extended in [23] using a forward reachability approach, which shows that any CBF is also a forward reachability solution within the maximum robust control invariant set. Compared to these results, we consider stability of the control invariant set implicitly defined through the predictive safety filter optimisation problem within a larger domain in a soft-constrained manner. Furthermore, it holds that if the prediction horizon of the predictive safety filter goes to infinity, the maximum control invariant set is also recovered, see, e.g., [24, Definition 2.9].
@@
Numerous works consider approximation methods to reduce the computational drawbacks of online predictive control methods. Works such as, e.g., [25, 26, 27] are able to guarantee closed-loop properties such as stability and robust constraint satisfaction with respect to input uncertainty. However, the safety filter policy is in general non-continuous and multi-valued, such that continuous approximations can lead to arbitrary approximation errors. Other works consider learning the optimal value function of an MPC, such as, e.g., [28, 29] to increase performance rather than to address general safety considerations. An overview for learning safety certificates is provided in [30] with a main focus on certificate synthesis for continuous-time systems. In [31], a control barrier function is learned from expert demonstrations with theoretical guarantees in case of Lipschitz continuity. These results could in principle be applicable for the proposed approximation scheme, however Lipschitz continuity of the predictive control barrier function is in general not guaranteed. In [32], a state-action control barrier function is approximated which makes use of contractivity of a safe set to ensure invariance of the resulting safe set for small enough approximation errors with a convex online optimisation problem. This results in the need to sample over the state-action space compared to sampling over the state space as proposed in this paper. Additionally, stability of the safe set itself is not considered in [32] and recursive feasibility only holds if the approximation error is small enough, whereas the proposed formulation ensures feasibility everywhere.
[…]
@@ 2 PRELIMINARIES
$x(k+1)=f(x(k),u(k))\;\;\forall k\in\mathbb{N},\$ (1)
with state $x\in\mathbb{R}^{n}$ and input $u\in\mathbb{R}^{m}$ and continuous dynamics $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$. The system is subject to state and input constraints
$\begin{split}x(k)&\in\mathcal{X}:=\{x\in\mathbb{R}^{n}\mid c_{x}(x)\leq 0\},\\ u(k)&\in\mathcal{U},\end{split}$ (2)
where the input constraints are considered to arise from, e.g., physical actuator limitations, while the state constraints arise from system specifications, such as, e.g., a limit on the velocity of a car on a highway lane. In practical predictive optimisation based methods, such constraints are typically formulated as soft constraints, see, e.g., [11], in order to preserve feasibility under unknown disturbances of the optimisation problems which are solved online. Accordingly, we define the soft-constrained set
@@
The constraint sets $\mathcal{U}$ and $\mathcal{X}(\xi)$ are compact for all $0\leq\xi<\infty$. Furthermore, it holds that $c_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n_{x}}$ is continuous.
The goal is to design a predictive safety filter for system (1). Given an initial state that lies outside the constraints $\mathcal{X}$, the system states should converge back into $\mathcal{X}$ and an initial state within a safe set should remain therein. In order to ensure such closed-loop stability and constraint satisfaction, we make use of the concept of control barrier functions. These functions typically define an invariant set, see, e.g. [33], as a zero sublevel set and provide stability guarantees within some domain with respect to this invariant set. Therefore, they are often considered to be an extension of a classical Lyapunov function, where the origin is stabilised. CBFs have been extensively studied in continuous-time dynamical systems, see, e.g., [5] for an overview, but have also been proposed in discrete-time in [16]. In this work, we consider two definitions for discrete-time CBFs introduced in [12] and [8]. These definitions differ in the properties the CBF needs to satisfy and we establish their relationship in Section 4 in order to analyse robust stability of the PCBF and the proposed algorithm.
@@
A function $h:\mathcal{D}\to\mathbb{R}$ is a discrete-time control barrier function with safe set $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}\subset\mathcal{D}\subset\mathbb{R}^{n}$ if the following conditions hold:
  1. (1)
@@
  3. (3)
    $\exists\Delta h:\mathcal{D}\to\mathbb{R}$, with $\Delta h$ continuous and $\Delta h(x)>0$ for all $x\in\mathcal{D}\setminus\mathcal{S}$ such that:
$\hskip 17.07182pt\forall x\in\mathcal{D}\setminus\mathcal{S}:$
$\inf_{u\in\mathcal{U}}\{h(f(x,u))\mid f(x,u)\in\mathcal{D}\}-h(x)\leq-\Delta h(x),$ (3a)
$\hskip 17.07182pt\forall x\in\mathcal{S}\>:\inf_{u\in\mathcal{U}}h(f(x,u))\leq 0.$ (3b)
This first definition requires that the CBF $h$ is continuous and defines a set $\mathcal{S}$ as a zero sublevel set which is control invariant through the condition (3b). Additionally, for any state outside the safe set, i.e., any $x\in\mathcal{D}\setminus\mathcal{S}$, there exists a continuous, positive decrease $\Delta h(x)$ of the CBF value. Inputs which satisfy the CBF conditions are considered safe to be applied to the given dynamical system.
@@
Given a CBF $h$ according to Definition 1, an input $u$ applied to system (1) at state $x$ is said to be safe, if $u\in\mathcal{U}_{h}(x)\subseteq\mathcal{U}$ with
@@
where $U_{1}(x)\coloneqq\{u\in\mathcal{U}\mid h(f(x,u))-h(x)\leq-\Delta h(x)\}$ and $U_{2}(x)\coloneqq\{u\in\mathcal{U}\mid h(f(x,u))\leq 0\}$.
@@
Let $\mathcal{S}$ and $\mathcal{D}$ be non-empty, compact and positively invariant sets for the system $x(k+1)=f(x(k),\kappa(x(k))$ with $\kappa:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ such that $\mathcal{S}\subset\mathcal{D}$. The set $\mathcal{S}$ is called an asymptotically stable set in $\mathcal{D}$, if for all $x(0)\in\mathcal{D}$, the following conditions hold:
@@
$\exists\bar{\delta}:$ $\;\|x(0)\|_{\mathcal{S}}<\bar{\delta}\Rightarrow\lim_{k\rightarrow\infty}x(k)\in\mathcal{S}$
Applying safe inputs to system (1) ensures that the set $\mathcal{S}$ is asymptotically stable in the domain $\mathcal{D}$ as given by the following result.
@@
Let $\mathcal{D}\subset\mathbb{R}^{n}$ be a non-empty and compact set. Consider a control barrier function $h:\mathcal{D}\rightarrow\mathbb{R}$ according to Definition 1 with $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}\subset\mathcal{D}$. If $\mathcal{D}$ is a forward invariant set for system (1) under $u(k)=\kappa(x(k))$ for all $\kappa:\mathcal{D}\rightarrow\mathcal{U}$ with $\kappa(x)\in\mathcal{U}_{h}(x)$, then it holds that
@@
The second definition of a control barrier function which we consider is given in [8, Appendix B.39] and makes use of class $\mathcal{K}$ comparator functions [^1] A function $\alpha:[0,c]\rightarrow\mathcal{R}$ is class $\mathcal{K}$ if it is continuous, strictly increasing, $\alpha(0)=0$..
@@
A function $h:\mathcal{D}\rightarrow\mathbb{R}$ is a discrete-time control barrier function with safe set $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)=0\}\subset\mathcal{D}$ if the following hold:
  1. (1)
@@
$\inf_{u\in\mathcal{U}}\{h(f(x,u))|\;f(x,u)\in\mathcal{D}\}-h(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}})$ (8)
A definition based on such comparator functions is advantageous due to the ease of extending stability guarantees if there are additional exogeneous disturbances acting on the system in (1). In Section 4, we will show that a CBF fulfilling Definition 1 implies existence of a CBF according to Definition 4. A similar relation of a continuous Lyapunov decrease implying the existence of a class $\mathcal{K}$ decrease is shown in [34], but uses class $\mathcal{K}_{\infty}$ bounds on the original Lyapunov function rather than the continuity of $h$ in Definition 1. The established relation between the two CBF definitions allows showing inherent robustness properties of the algorithm in [12]. Furthermore, it enables the input-to-state stability guarantees of the approximate predictive CBF approach proposed in this paper.
@@
3 PREDICTIVE CONTROL BARRIER FUNCTIONS
In this section, we review the concept of predictive safety filters, as proposed originally in [7] and adapted in [12], and how they can be related to the concept of a control barrier function (CBF). At every time step, a controller proposes a potentially unsafe control action, denoted as $u_{\mathrm{p}}(k)$. Such a controller could be optimising a primary control objective, such as a robot operating in a cluttered environment or even a human operating a car. The aim of the safety filter is to certify, whether the proposed input is safe to apply to the system in terms of the state and input constraints (2), and modify it if necessary. In order to guarantee constraint satisfaction for all times, the predictive safety filter makes use of model predictive control techniques. The predictive safety filter problem proposed in [12] which is solved online at every time step is given as
$\min_{u_{i|k}}\;\;$ $\lvert\lvert{u_{0|k}-u_{\mathrm{p}}(k)}\rvert\rvert$ (9a)

### v2

@@ Abstract
We propose integrating an approximation of a predictive control barrier function (PCBF) in a safety filter framework, resulting in a prediction horizon independent formulation. The PCBF is defined through the value function of an optimal control problem and ensures invariance as well as stability of a safe set within a larger domain of attraction. We provide a theoretical analysis of the proposed algorithm, establishing input-to-state stability of the safe set with respect to approximation errors as well as exogenous disturbances. Furthermore, we propose a continuous extension of the PCBF within the safe set, reducing the impact of learning errors on filter interventions. We demonstrate the stability properties and computational advantages of the proposed algorithm on a linear system example and its application as a safety filter for miniature race cars in simulation.
@@ 1 INTRODUCTION
Safety filters provide a modular mechanism which enables certifying and, if necessary, modifying any control input such that safety in terms of constraint satisfaction can be guaranteed. The filters are designed to be agnostic with respect to the main control performance objective and therefore aim to minimise the distance between the proposed input and the filtered input which is applied to the system. Thereby, safety filters are suitable tools to enhance the safety properties of various systems. The filters can be used to augment human agents operating systems, such as, e.g., automotive vehicles as demonstrated on miniature race cars in [1], and have also demonstrated their effectiveness in complex robotics applications, such as, e.g., walking on stepping stones with a legged robot in [2].
@@
The core principle of safety filters consists in obtaining a control invariant set within the state constraints, such that inputs applied to the system guarantee that the system state remains within this invariant set. In order to obtain such a control invariant set, three main branches have been pursued, as discussed extensively in [3]. Hamilton-Jacobi reachability analysis, as discussed in [4], allows computing the maximal control invariant subset to constraints, however the method suffers from poor computational scalability with respect to increasing state dimension. Control barrier functions (CBFs) aim to alleviate the computational restrictions, by explicitly computing, potentially smaller, control invariant sets. An overview is provided in [5]. CBFs are obtained either through either expert knowledge or more scalable computational methods compared to the Hamilton-Jacobi reachability, such as semi-definite programming or sum-of-squares optimisation, such as in [6]. Finally, predictive safety filters have been proposed in [7]. They employ model predictive control (MPC) techniques in order to implicitly define a control invariant set through an online predictive optimisation problem. A detailed overview of such MPC techniques can be found in [8]. These predictive methods make use of discrete-time dynamics formulations and allow to easily provide recursive feasibility guarantees. They enable extensions to, e.g., robust treatments of disturbances, stochastic disturbances in [9] or using online data for uncertain systems to improve safety filter performance in [10], due to the well developed MPC theory.
@@
In practical systems, considering hard constraints in predictive control methods often leads to infeasibility due to unmodelled dynamics or sensor noise. Therefore, the state constraints on the system are often softened, as proposed in [11] and used in [1], by introducing slack variables. However, this generally leads to a loss of theoretical properties as soon as the system violates the constraints. In [12], a constraint softening mechanism is proposed, which is able to provide theoretical guarantees when constraints are violated. Connections of this mechanism to CBFs are investigated and it is shown that by minimising a function which depends on the slack variables, a control barrier function is implicitly defined. This result implies that the implicitly defined safe set of a predictive safety filter is stable within a larger domain of attraction, such that the system is able to recover from unexpected disturbances or model mismatch. The main drawback of the method in [12] is the need to solve two predictive optimisation problems in order to guarantee stability, which results in a computationally demanding algorithm.
Contributions. In this work, we leverage an explicit approximation of the PCBF in a safety filter formulation. This formulation requires predicting only a single time step ahead, allowing for efficient online computations. The contributions compared to initial results in [13] are as follows:
  1. 1.
    We extend initial input-to-state stability (ISS) analysis in [13], with respect to approximation errors, to the case of additional exogenous disturbances. This is achieved through a relation between control barrier function (CBF) definitions based on continuity and class $\mathcal{K}$ comparison functions, which enables a general robust set stability analysis. The provided analysis directly implies ISS of the PCBF algorithm in [12].
  2. 2.
    We propose a novel continuous extension of the PCBF, leading to a significant reduction in safety filter interventions when the system is safe compared to [13].
  3. 3.
    We introduce a hyperparameter that reduces the required CBF decrease in closed-loop compared to [13] and [12], while preserving stability guarantees.
  4. 4.
    We show how the explicit approximation can be used in order to certify inputs using a single function evaluation, resulting in large computational speedups compared to [13, 12] when safe inputs are proposed.
@@
Related Work. Safety filters have been extensively researched and a review of state-of-the-art methods can be found in [3]. The first main branch of research consists of control barrier function based safety filters, originally proposed in [15] for continuous time systems. The work in [16] proposed discrete-time control barrier functions. However, the proposed formulation uses a barrier function which approaches infinity as the state approaches the boundary of the safe set, which is not suitable for investigating the stability of the safe set within a larger domain. In [17], a backup CBF formulation is proposed, which allows using a backup controller to enhance an invariant safe set, by computing gradients of a CBF on the closed-loop trajectory of the backup controller. Thereby, any state which reaches the safe set under the backup controller is still considered safe, and the safe set is effectively augmented. The predictive safety filters in [18] and [12] show relations to this method, where the optimiser is free to choose all control inputs in the prediction horizon, rather than selecting them with a prior fixed policy. In [19], the predictive safety filter is augmented with a stability guarantees of the origin using a Lyapunov decrease constraint. Compared to these works, we combine the advantages of the existence of inputs which stabilise a safe set and render it invariant, while using a single CBF decrease constraint rather than a full prediction horizon, reducing the computational effort. Due to the discrete-time formulation of the predictive controllers, the resulting CBF optimisation problem is a nonlinear program rather than a quadratic program in [17], however the prediction horizon is not actually included in the online optimisation, effectively reducing the number of constraints. The proposed formulation in [20] uses a look-ahead to determine time instances, where given a path of the system, a CBF constraint is violated. Thereby, the control actions are modified such that safety can be guaranteed. However, this method does not consider the case when the input is constrained. Finally, [21] proposes a reachability-based safety filter formulation in continuous time, which makes use of a finite lookahead time and establishes connections to model predictive safety filter methods. However, the method uses numerical approximations of the gradients of the CBF through sampling to impose the CBF decrease and it is unclear how approximation errors can impact the closed-loop performance. In comparison, we provide a rigorous analysis of how approximation errors of the learned CBF impact the closed-loop behaviour of the system.
@@
Another main branch of research for safety filters is Hamilton-Jacobi reachability analysis, see, e.g., [4]. In [22], connections between viscosity solution of the Hamilton-Jacobi-Isaacs variational inequality, retrieving the maximum robust control invariant subset to given constraints and CBFs are established and it is shown how this solution can be used in a quadratic program. This result is extended in [23] using a forward reachability approach, which shows that any CBF is also a forward reachability solution within the maximum robust control invariant set. Compared to these results, we consider stability of the control invariant set implicitly defined through the predictive safety filter optimisation problem within a larger domain in a soft-constrained manner. Furthermore, if the prediction horizon of the predictive safety filter goes to infinity, the maximum control invariant set can be recovered, see, e.g., [24, Definition 2.9].
@@
Numerous works consider approximation methods to reduce the computational drawbacks of online predictive control methods. Works such as, e.g., [25, 26, 27] are able to guarantee closed-loop properties such as stability and robust constraint satisfaction with respect to input uncertainty. However, the safety filter policy is in general non-continuous and multi-valued, such that continuous approximations can lead to arbitrary approximation errors. Other works consider learning the optimal value function of an MPC, such as, e.g., [28, 29] to increase performance rather than to address general safety considerations. An overview for learning safety certificates is provided in [30] with a main focus on certificate synthesis for continuous-time systems. In [31], a control barrier function is learned from expert demonstrations with theoretical guarantees in case of Lipschitz continuity. These results could in principle be applicable for the proposed approximation scheme, however Lipschitz continuity of the predictive control barrier function is in general not guaranteed. Additionally, the required optimization problem is computationally complex, especially for neural network approximations. A discussion on loss function design for neural network approximations is provided in [49]. In comparison, we provide guarantees on closed-loop system behaviour even if approximation errors are still present in the learned CBF and additionally consider the effect of exogenous disturbances. In [32], a state-action control barrier function is approximated which makes use of contractivity of a safe set to ensure invariance of the resulting safe set for small enough approximation errors with a convex online optimisation problem. This results in the need to sample over the state-action space compared to sampling over the state space as proposed in this paper. Additionally, stability of the safe set itself is not considered in [32] and recursive feasibility only holds if the approximation error is small enough, whereas the proposed formulation ensures feasibility everywhere.
Structure. We discuss preliminaries in Section 2 and summarize the PCBF algorithm in [12], which is the basis for the approximation used in this paper, in Section 3. In Section 4, we detail the proposed algorithm and the continuous extension of the PCBF, and in Section 5, we provide a relationship between two types of CBF definitions as well as the theoretical closed-loop analysis of the proposed algorithm. In Section 6, we validate the approach in two numerical examples.
Notation. Given a vector $a\in\mathbb{R}^{n}$ and a scalar $b\in\mathbb{R}\cup\{{\infty}\}$, an inequality $a<b$ denotes an element-wise inequality of each element of $a$, i.e. $a]_{i}<b$ for all $i=1,\dots,n$. The distance of a vector $x$ to a set $\mathcal{A}$ is given by $\|x\|_{\mathcal{A}}:=\inf_{y\in\mathcal{A}}\|x-y\|$. A vector of appropriate dimensions with all entries equal to 1 is denoted as $\mathbf{1}$.
@@ 2 PRELIMINARIES
$x(k+1)=f(x(k),u(k),w(k))\;\;\forall k\in\mathbb{N},$ (1)
with state $x\in\mathbb{R}^{n}$, input $u\in\mathbb{R}^{m}$, exogenous disturbance $w(k)\in\mathbb{R}^{p}$ and continuous dynamics $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{n}$. The system is subject to state and input constraints
$x(k)\in\mathcal{X}:=\{x\in\mathbb{R}^{n}\mid c_{x}(x)\leq 0\},\;\;u(k)\in\mathcal{U},$ (2)
with $c_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n_{x}}$. The input constraints are considered to arise from, e.g., physical actuator limitations, while the state constraints arise from system specifications, such as, e.g., a limit on the velocity of a car on a highway lane. In practical predictive optimisation based methods, such constraints are typically formulated as soft constraints, see, e.g., [11], in order to preserve feasibility under unknown disturbances of the optimisation problems which are solved online. Accordingly, we define the soft-constrained set
@@
The sets $\mathcal{U}$ and $\mathcal{X}(\xi)$ are compact for all $0\leq\xi<\infty$ and $c_{x}$ is continuous.
The goal is to design a predictive safety filter for system (1). If an initial state violates the constraints $\mathcal{X}$, the system states should converge back into $\mathcal{X}$ and an initial state within a safe set should remain therein. In order to ensure such closed-loop stability and constraint satisfaction, we make use of the concept of control barrier functions. These functions typically define an invariant set, see, e.g. [52], as a zero sublevel set and provide stability guarantees within some domain with respect to this invariant set. Therefore, they are often considered to be an extension of a classical Lyapunov function, where the origin is stabilised. CBFs have been extensively studied in continuous-time dynamical systems, see, e.g., [5] for an overview, but have also been proposed in discrete-time in [16]. In this work, we consider two definitions for discrete-time CBFs introduced in [12] and [8]. These definitions differ in the properties the CBF needs to satisfy and we establish their relationship in Section 5 in order to analyse robust stability of the proposed algorithm.
@@
A function $h:\mathcal{D}\to\mathbb{R}$ is a discrete-time CBF with safe set $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}\subset\mathcal{D}\subset\mathbb{R}^{n}$ if the following conditions hold:
  1. 1)
@@
  3. 3)
    $\exists\Delta h:\mathcal{D}\to\mathbb{R}$, with $\Delta h$ continuous and $\Delta h(x)>0$ for all $x\in\mathcal{D}\setminus\mathcal{S}$ such that
$\hskip 19.91684pt\forall x\in\mathcal{D}\setminus\mathcal{S}:$
$\inf_{u\in\mathcal{U}}\{h(f(x,u,0))\mid f(x,u,0)\in\mathcal{D}\}\!-\!h(x)\!\leq\!-\Delta h(x),$ (3a)
$\hskip 20.48601pt\forall x\in\mathcal{S}\>:\inf_{u\in\mathcal{U}}h(f(x,u,0))\leq 0.$ (3b)
This first definition requires that the CBF $h$ is continuous and defines a set $\mathcal{S}$ as a zero sublevel set which is control invariant through the condition (3b). Additionally, for any state outside the safe set, i.e., any $x\in\mathcal{D}\setminus\mathcal{S}$, there exists a continuous, positive decrease $\Delta h(x)$ of the CBF value. We note that compared to other CBF definitions such as in [16], we consider the 0-sublevel set as the safe set rather than the superlevel set. If inputs that satisfy (3) are selected, it was shown in [12] that the set $\mathcal{S}$ is asymptotically stable in $\mathcal{D}$ for the nominal closed-loop system by leveraging continuity-based arguments. Inputs which satisfy the CBF conditions are considered safe to be applied to the given dynamical system.
@@
Given a CBF $h$ according to Definition 1, an input $u$ applied to system (1) at state $x$ is said to be safe for the nominal system, if $u\in\mathcal{U}_{h}(x)\subseteq\mathcal{U}$ with
@@
where $U_{1}(x)\coloneqq\{u\in\mathcal{U}\mid h(f(x,u,0))-h(x)\leq-\Delta h(x)\}$ and $U_{2}(x)\coloneqq\{u\in\mathcal{U}\mid h(f(x,u,0))\leq 0\}$.
@@
Let $\mathcal{S}$ and $\mathcal{D}$ be non-empty, compact and positively invariant sets for the system $x(k+1)=f(x(k),\kappa(x(k)),0)$ with $\kappa:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ such that $\mathcal{S}\subset\mathcal{D}$. The set $\mathcal{S}$ is called an asymptotically stable set in $\mathcal{D}$, if for all $x(0)\in\mathcal{D}$, the following conditions hold:
@@
$\exists\bar{\delta}>0:$ $\;\|x(0)\|_{\mathcal{S}}<\bar{\delta}\Rightarrow\lim_{k\rightarrow\infty}x(k)\in\mathcal{S}$
Applying safe inputs to system (1) ensures that the set $\mathcal{S}$ is asymptotically stable in the domain $\mathcal{D}$ for the nominal system as given by the following result.
@@
Let $\mathcal{D}\subset\mathbb{R}^{n}$ be a non-empty and compact set. Consider a control barrier function $h:\mathcal{D}\rightarrow\mathbb{R}$ according to Definition 1 with $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}\subset\mathcal{D}$. If $\mathcal{D}$ is a forward invariant set for system $x(k+1)=f(x(k),u(k),0)$ under $u(k)=\kappa(x(k))$ for all $\kappa:\mathcal{D}\rightarrow\mathcal{U}$ with $\kappa(x)\in\mathcal{U}_{h}(x)$, then it holds that
@@
The second definition of a control barrier function which we consider is given in [8, Appendix B.39] and makes use of class $\mathcal{K}$ comparison functions [^1] A function $\alpha:[0,c]\rightarrow\mathcal{R}$ is class $\mathcal{K}$ if it is continuous, strictly increasing, $\alpha(0)=0$..
@@
A function $h:\mathcal{D}\rightarrow\mathbb{R}$ is a discrete-time CBF with safe set $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}\subset\mathcal{D}$ if the following hold:
  1. 1)
@@
$\inf_{u\in\mathcal{U}}\{h(f(x,u,0))\mid f(x,u,0)\in\mathcal{D}\}-h(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}}).$
@@
A definition based on such comparison functions is advantageous due to the ease of extending stability guarantees if there are additional exogenous disturbances acting on the system in (1). In Section 5, we will show that a CBF fulfilling Definition 1 implies existence of a CBF according to Definition 4. A similar relation of a continuous Lyapunov decrease implying the existence of a class $\mathcal{K}$ decrease is shown in [34], but uses class $\mathcal{K}_{\infty}$ bounds on the original Lyapunov function rather than the continuity of $h$ in Definition 1. The established relation between the two CBF definitions allows showing the input-to-state stability guarantees of the approximate predictive CBF approach proposed in this paper.
3 PRELIMINARIES ON PREDICTIVE CONTROL BARRIER FUNCTIONS
In this section, we review the concept of predictive safety filters, as proposed originally in [7] and adapted in [12] for system recovery, and how they can be related to the concept of CBFs. At every time step, a controller proposes a potentially unsafe control action, denoted as $u_{\mathrm{p}}(k)$. Such a controller could be optimising a primary control objective, such as a robot operating in a cluttered environment or even a human operating a car. The aim of the safety filter is to certify whether the proposed input is safe to apply to the system in terms of the state and input constraints (2), and modify it, if necessary. In order to guarantee constraint satisfaction for all times, the predictive safety filter makes use of model predictive control techniques. The predictive safety filter problem proposed in [12] which is solved online at every time step is given as
$\min_{x_{i|k},u_{i|k}}\;\;$ $\lvert\lvert{u_{0|k}-u_{\mathrm{p}}(k)}\rvert\rvert$ (8a)